OpenStudy (anonymous):

2x 4 ---- + ---- = 5 solve for x 7 5

OpenStudy (anonymous):

should be 2.8

OpenStudy (anonymous):

wait i didnt see the =5

OpenStudy (anonymous):

\[\frac{ 2x }{7 }+\frac{ 4 }{5}=5\] Subtract \[\frac{ 2x }{7 }+\frac{ 4 }{5}-\frac{ 4 }{5}=5-\frac{ 4 }{5}\] = \[\frac{ 2x }{7 }=\frac{ 25 }{ 5 }-\frac{ 4 }{ 5 }\] = \[\frac{ 2x }{7 }=\frac{ 21 }{ 5 }\] Multiply by 7 \[\frac{ 2x }{7 }*7=\frac{ 21 }{ 5 }*7\] = \[2x=\frac{ 147 }{5 }\] Divide by 2 = \[2x*2=29.4*2\] = \[x=58.8\]

OpenStudy (anonymous):

I multiplied sorry \[x=\frac{ 147 }{10 }\] 14.7